This problem involves solving demand and supply equations together to determine price and quantity. QD = -2P +20 QS = 2P – 4


1
Expert's answer
2021-09-09T17:56:27-0400

"Qd=Qs\\-2P+20=2P-4\\4P=24\\P=6\\Q=2(6)-4\\=12-4\\Q=8"